OpenStudy (josee):

Find the value of f'(a) when f(x) = sqrt(5x + 6)

OpenStudy (josee):

Omg so many ppl are typing :3 i feel so loved ! lol

OpenStudy (anonymous):

\[\frac{1}{\sqrt{5a+6}}*\frac{1}{2}*5\]

OpenStudy (anonymous):

f'(a)=1/2(5(a)+6)^-1/2(5)

OpenStudy (lalaly):

\[f(a)=\sqrt{5a+6} =(5a+6)^{\frac{1}{2}}\]\[f'(a) = \frac{5}{2}(5a+6)^{\frac{-1}{2}}\]\[=\frac{5}{2\sqrt{5a+6}}\]

OpenStudy (josee):

Thank you lalaly!:D